a rectangles lengeth is five inches less than four times its width. if the perimeter is fourty inches what are the dimensions of the rectangle

Found 2 solutions by oberobic, unlockmath:
Answer by oberobic(2304)

L = length
W = width
P = 2L + 2W = perimeter
.
P = 40 :: given
.
L = 4W -5 :: length is 5 less than 4*Width
.
Substitute and solve for W.
.
2L + 2W = P
2(4W-5) + 2W= 40
8W -10 + 2W = 40
10W -10 = 40
10W = 50
W = 5
.
L = 4W-5
L = 4(5) -5
L = 20 - 5
L = 15
.
Checking the perimeter will tell if this is the right answer
2L + 2W = 40??
.
2(15) = 30
2(5) = 10
30+10 = 40
Correct.
.
Answer:
The dimensions of the rectangle are length = 15 and width = 5.
.
Done.

Answer by unlockmath(1599)

Hello,
Let the width be x and the length will be 4x-5. Now we can set up our equation as:
x+x+(4x-5)+(4x-5)=40 Combine like terms to get:
10x-10=40 Add 10 to both sides and divide by 10 to get:
x=5
There we go the width is 5 inches. I'll let you figure out the width.
(Note: Lenegth is spelled length and fourty is spelled forty)
RJ
Check out a book I wrote:
www.math-unlock.com